OpenStudy (anonymous):

Perform the indicated operation. 2/3 - 5/8

jimthompson5910 (jim_thompson5910):

The denominators are not the same, so you can't subtract until they are

jimthompson5910 (jim_thompson5910):

What is the LCD of 3 and 8?

OpenStudy (anonymous):

LCD?

jimthompson5910 (jim_thompson5910):

or LCM of 3 and 8

jimthompson5910 (jim_thompson5910):

LCD = lowest common denominator LCM = lowest common multiple

OpenStudy (anonymous):

Would I have to cross multiply?

jimthompson5910 (jim_thompson5910):

that only applies to equations, but I think I know what you mean

OpenStudy (anonymous):

1\24 is the answer

jimthompson5910 (jim_thompson5910):

\[\Large \frac{2}{3} - \frac{5}{8}\] \[\Large \frac{2*8}{3*8} - \frac{5*3}{8*3}\] \[\Large \frac{16}{24} - \frac{15}{24}\] Was that what you were thinking?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Yes Jim that is what I was thinking. Now would I need to subtract just the top?

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

So it would be 1/24

jimthompson5910 (jim_thompson5910):

you nailed it

jimthompson5910 (jim_thompson5910):

hopefully you see how I got all that?

OpenStudy (anonymous):

I did yes, and It's helping me on my other questions as well :)

jimthompson5910 (jim_thompson5910):

alright, that's great